# A geometric approach toconjugation-invariant random permutations

### v3

@@ 5 Proofs of the results on Robinson–Schensted shapes
Proof of Theorem 2.7.
Fix $r\geq 0$. Let $\mathcal{P}$ be a geometric construction of $\tau_{n}$. The reason we can apply the results of [Sjö23] is that $\mathcal{P}$ is locally, outside the diagonal, close to a homogeneous Poisson point process with intensity $n$. More precisely, let $I,J$ be essentially disjoint intervals of $[0,1]$ and write $\bigtriangleup$ for the symmetric difference of sets. Thanks to 3.2 and 3.3, one can construct a homogeneous Poisson point process with intensity $n$ on $I\times J$, say $\sigma_{n}$, such that ${\left|\sigma_{n}\bigtriangleup\mathcal{P}\right|}/{n}\to 0$ as $n\to\infty$ almost surely.
@@
Figure 4: The bands defined in the proof of Theorem 2.7. On each $A_{i}$ and each $B_{i}$, $\mathcal{P}$ is almost distributed like a homogeneous Poisson point process.
@@
for $i\in[K]$, see Figure 4. These bands each fit in the setting of 3.3, and they satisfy $\mathrm{Leb}(C)=\frac{1}{K+1}$ where $C:=(0,1)^{2}\setminus\left(\bigcup_{i}\bar{A}_{i}\cup\bar{B}_{i}\right)$. Moreover, using 3.2 and 3.3 as above, we can construct a point process $\sigma_{n}$ on $(0,1)^{2}$ such that:
@@
For each $i\in[K]$, we can apply [Sjö23, Lemma 8.2] to $\sigma_{n}$ on $A_{i}$. Hence for any $u\in\mathcal{U}(A_{i})$, there exists $\delta_{u,A_{i}}>0$ such that:
$\sup_{w\in\mathcal{U}(A_{i}),\,\lVert w-u\rVert_{1,A_{i}}<\delta_{u,A_{i}}}\left|D(w\sqrt{n})\cap\sigma_{n}\cap A_{i}\right|/n<F_{A_{i}}(u)+\epsilon$
@@
$\sup_{w\in\mathcal{U}(C)}\left|D(w\sqrt{n})\cap\sigma_{n}\cap C\right|/n<\left|\sigma_{n}\cap C\right|/n=\left|\mathcal{P}\cap C\right|/n.$
Since the property of being doubly increasing is stable by restriction, for any $u\in\mathcal{U}\left((0,1)^{2}\right)$ we can write
$F_{(0,1)^{2}}(u)=F_{C}(u)+\sum_{i=1}^{K}(F_{A_{i}}(u)+F_{B_{i}}(u))\geq\sum_{i=1}^{K}(F_{A_{i}}(u)+F_{B_{i}}(u)).$
By summing the previous inequalities we deduce that for any $u\in\mathcal{U}\left((0,1)^{2}\right)$, with $\delta_{u}:=\min_{1\leq i\leq K}\delta_{u,A_{i}}\wedge\delta_{u,B_{i}}$:
$\sup_{w\in\mathcal{U}\left((0,1)^{2}\right),\,\lVert w-u\rVert_{1,(0,1)^{2}}<\delta_{u}}\left|D(w\sqrt{n})\cap\sigma_{n}\right|/n<F_{(0,1)^{2}}(u)+2K\epsilon+\left|\mathcal{P}\cap C\right|/n$
@@
$\sup_{w\in\mathcal{U}\left((0,1)^{2}\right),\,\lVert w-u\rVert_{1,(0,1)^{2}}<\delta_{u}}\left|D(w\sqrt{n})\cap\mathcal{P}\right|/n<F_{(0,1)^{2}}(u)+2K\epsilon+\left|\mathcal{P}\cap C\right|/n+\left|\sigma_{n}\bigtriangleup\mathcal{P}\right|/n$
w.h.p. as $n\to\infty$. Recall that $\epsilon=1/K^{2}$ and a.s. $\left|\sigma_{n}\bigtriangleup\mathcal{P}\right|/n\to 0$ as $n\to\infty$. Moreover 3.2 implies that a.s. $\left|\mathcal{P}\cap C\right|/n\to\mathrm{Leb}(C)=1/(K+1)$ as $n\to\infty$. Hence for any $u\in\mathcal{U}\left((0,1)^{2}\right)$:
$\sup_{w\in\mathcal{U}\left((0,1)^{2}\right),\,\lVert w-u\rVert_{1,(0,1)^{2}}<\delta_{u}}\left|D(w\sqrt{n})\cap\mathcal{P}\right|/n<F_{(0,1)^{2}}(u)+4/K$
w.h.p. as $n\to\infty$. According to [Sjö23, Proposition 9.2], $\mathcal{U}_{0,r}\left((0,1)^{2}\right)$ is compact for the $L^{1}$ metric. Thus its open cover
$\mathcal{U}_{0,r}\left((0,1)^{2}\right)=\bigcup_{u\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)}\left\{w\,:\,\lVert w-u\rVert_{1,(0,1)^{2}}<\delta_{u}\right\}$
admits a finite subcover, and we can conclude that with high probability as $n\to\infty$:
for $\text{for any }w\in\mathcal{U}_{0,r}\left((0,1)^{2}\right),\quad\left|D(w\sqrt{n})\cap\mathcal{P}\right|/n<\sup_{u\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)}F_{(0,1)^{2}}(u)+4/K.$
Using Equations 14 and 13, this yields
$\frac{1}{n}\mathrm{LDS}_{r\sqrt{n}}(\mathcal{P})<F_{\mathrm{LSKV}}(r)+4/K$ (15)
@@
The lower bound requires more work, and we need to delve deeper into the proofs of [Sjö23]. The reader is thus kindly invited to follow along with his own copy of [Sjö23]. Our proof hinges on the notation of [Sjö23, Definition 7.3], the results [Sjö23, Lemmas 4.9, 6.3, 7.7, 9.1], and the method behind [Sjö23, Lemma 10.1].
Recall (14) and fix $u\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)$ such that
$F_{(0,1)^{2}}(u)>F_{\mathrm{LSKV}}(r)-1/K.$ (16)
Our goal is to construct a large enough $r\sqrt{n}$-decreasing subset of $\mathcal{P}$, i.e. by (13) to construct $w_{n}\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)$ such that $\left|D(w_{n}\sqrt{n})\cap\mathcal{P}\right|/n\geq F_{(0,1)^{2}}(u)-\epsilon^{\prime}$ for some small $\epsilon^{\prime}$ and large enough $n$. Unfortunately we can not simply apply [Sjö23, Lemma 10.1] to each $A_{i}$ and $B_{i}$, as combining the resulting local, doubly increasing, functions into a single global, doubly increasing, function on $(0,1)^{2}$ would be difficult. Instead we rewrite the proof of this result in our framework: everything works the same, with the exception that we need to work on each $A_{i}$ and $B_{i}$ individually.
First apply [Sjö23, Lemma 7.7] to each domain $A_{i}\cup B_{i}$ with constant density $\rho=1$: since the density is constant, some items of this lemma become vacuous, and we only rewrite the relevant ones. Recall that we use the notation of [Sjö23, Definition 7.3]. For any $\iota\in(0,1/2)$ there exists a measurable subset $S_{\iota}^{i}\subseteq A_{i}\cup B_{i}$ on which $u$ is differentiable, and a finite collection $\mathfrak{P}_{\iota}^{i}$ of disjoint $(u,\iota)$-parallelograms, such that:
  • (a)
    $S_{\iota}^{i}\subseteq\bigcup_{P\in\mathfrak{P}_{\iota}^{i}}P$;
  • (b)
    $\partial_{x}u$ and $\partial_{y}u$ are bounded on $\bigcup_{0<\iota<1/2}S_{\iota}^{i}$;
  • (c)
    $\mathrm{Leb}(A_{i}\cup B_{i}\setminus S_{\iota}^{i})<\epsilon+o_{\iota}(1)$;
and for each $P\in\mathfrak{P}_{\iota}^{i}$:
  • (a’)
    $P\subseteq A_{i}\cup B_{i}$, with center $(x_{P},y_{P})\in S_{\iota}^{i}$;
  • (f)
    if we denote by $u_{x}^{P},u_{y}^{P}$ the partial derivatives of $u$ at $(x_{P},y_{P})$, then for all $(x,y)\in P$:
  • (g)
    $\left|\frac{1}{\mathrm{Leb}(P)}F_{P}(u)-L(1,u_{x}^{P}u_{y}^{P})\right|<\iota$.
Then for each $\iota$, set $S_{\iota}:=\bigcup_{i\in[K]}S_{\iota}^{i}$ and $\mathfrak{P}_{\iota}:=\bigcup_{i\in[K]}\mathfrak{P}_{\iota}^{i}$. Note that, given our definition, any parallelogram contained in $A_{i}\cup B_{i}$ is either wholly contained in $A_{i}$, or wholly contained in $B_{i}$. Now let us follow the proof of [Sjö23, Lemma 10.1].
@@
Fix $\iota\in(0,1/2)$. For each $P\in\mathfrak{P}_{\iota}$, let $P^{\prime}$ denote this parallelogram shrunk by a factor $1\!-\!2\iota$ in width and height. Denote by $\widetilde{\mathfrak{P}}_{\iota}$ the set of well-behaved parallelograms in $\mathfrak{P}_{\iota}$, i.e. for which we have $u_{x}^{P}\geq\iota^{3}$ and $u_{y}^{P}\geq\iota^{3}$. For each $P\in\mathfrak{P}_{\iota}$, let $\bar{u}_{P}\in\mathcal{U}(P)$ be defined by
for $\text{for all }(x,y)\in P,\quad\bar{u}_{P}(x,y):=u(x_{P},y_{P})+(x-x_{P})u_{x}^{P}+(y-y_{P})u_{y}^{P}$
@@
if $P\in\widetilde{\mathfrak{P}}_{\iota}$, and by $\bar{u}_{P}:=u$ on $P$ if $P\in\mathfrak{P}_{\iota}\setminus\widetilde{\mathfrak{P}}_{\iota}$. Let $r_{P}:=\mathrm{diam}\,\bar{u}_{P}(P^{\prime})$. Since $\sigma_{n}$ is distributed like a homogeneous Poisson point process with intensity $n$ on each $A_{i}$ and each $B_{i}$, we can apply [Sjö23, Lemma 6.3] to $\bar{u}_{P}$ and obtain the following analog of [Sjö23, Equation (18)]. For any $P\in\widetilde{\mathfrak{P}}_{\iota}$, w.h.p. as $n\to\infty$ there exists $w_{P}^{(n)}\in\mathcal{U}_{u(x_{P},y_{P})-\frac{r_{P}}{2},r_{P}}(P^{\prime})$ such that
@@
$\frac{1}{\mathrm{Leb}(P^{\prime})}\left|D\left(w_{P}^{(n)}\sqrt{n}\right)\cap\sigma_{n}\cap P^{\prime}\right|/n\geq L(1,u_{x}^{P}u_{y}^{P})-o_{\iota}(1)$ (17)
where $o_{\iota}(1)\to 0$ as $\iota\to 0$, uniformly in $P$ and $n$. By [Sjö23, Lemma 4.9] and item (b), the function $L$ is uniformly continuous on the bounded set $\left\{\left(1,\partial_{x}u(x,y)\,\partial_{y}u(x,y)\big{.}\right)\,:\,(x,y)\in\bigcup_{\iota}S_{\iota}\right\}$ and satisfies $L(1,0)=0$. By definition, for any $P\in\mathfrak{P}_{\iota}\setminus\widetilde{\mathfrak{P}}_{\iota}$ we have $\min\left(u_{x}^{P},u_{y}^{P}\right)<\iota^{3}$. Therefore for any $P\in\mathfrak{P}_{\iota}\setminus\widetilde{\mathfrak{P}}_{\iota}$ and any $w_{P}^{(n)}\in\mathcal{U}(P^{\prime})$, Equation 17 still holds.
@@
Under the event that $w_{P}^{(n)}$ exists for all $P\in\widetilde{\mathfrak{P}}_{\iota}$, define a function $w_{n}$ on $\bigcup_{P\in\widetilde{\mathfrak{P}}_{\iota}}P^{\prime}\,\cup\,\left((0,1)^{2}\setminus\bigcup_{P\in\widetilde{\mathfrak{P}}_{\iota}}P\right)$ as follows: for any $P\in\widetilde{\mathfrak{P}}_{\iota}$, $w_{n}$ equals $w_{P}^{(n)}$ on $P^{\prime}$, and $w_{n}$ equals $u$ on $(0,1)^{2}\setminus\bigcup_{P\in\widetilde{\mathfrak{P}}_{\iota}}P$. As in the proof of [Sjö23, Lemma 10.1], $w_{n}$ is doubly increasing on this domain (this mainly uses [Sjö23, Definition 7.3], the definition of $P^{\prime}$, and item (f)). Then by [Sjö23, Lemma 9.1], this function can be extended to $w_{n}\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)$. Therefore, using (17), we can write w.h.p. as $n\to\infty$:
@@
$\left|D(w_{n}\sqrt{n})\cap\sigma_{n}\right|/n\geq\sum_{P\in\mathfrak{P}_{\iota}}\left|D(w_{n}\sqrt{n})\cap\sigma_{n}\cap P^{\prime}\right|/n\geq\sum_{P\in\mathfrak{P}_{\iota}}\mathrm{Leb}(P^{\prime})\left(L(1,u_{x}^{P}u_{y}^{P})-o_{\iota}(1)\right)$ (18)
where $o_{\iota}(1)$ is independent of $P$ and $n$. By item (g):
$\sum_{P\in\mathfrak{P}_{\iota}}\mathrm{Leb}(P^{\prime})L(1,u_{x}^{P}u_{y}^{P})\geq(1\!-\!2\iota)^{2}\sum_{P\in\mathfrak{P}_{\iota}}\mathrm{Leb}(P)L(1,u_{x}^{P}u_{y}^{P})\geq(1\!-\!2\iota)^{2}F_{\bigcup_{P\in\mathfrak{P}_{\iota}}\!\!P}(u)-o_{\iota}(1).$ (19)
By [Sjö23, Lemma 4.9], $0\leq L(1,\partial_{x}u\,\partial_{y}u)\leq 1$, and by items (a) and (c):
$\mathrm{Leb}\left(\big{.}\bigcup_{P\in\mathfrak{P}_{\iota}}P\right)\geq\mathrm{Leb}\left(\bigcup_{i\in[K]}A_{i}\cup B_{i}\right)-K\epsilon-o_{\iota}(1)=1-2/K-o_{\iota}(1).$
@@
This implies $F_{\bigcup_{P\in\mathfrak{P}_{\iota}}\!\!P}(u)\geq F_{(0,1)^{2}}(u)-2/K-o_{\iota}(1)$. Along with (18) and (19), this yields:
$\left|D(w_{n}\sqrt{n})\cap\sigma_{n}\right|/n\geq F_{(0,1)^{2}}(u)-2/K-o_{\iota}(1)$
@@
w.h.p. as $n\to\infty$, where $o_{\iota}(1)$ is uniform in $n$. Take $\iota$ such that this $o_{\iota}(1)$ is less that $1/K$. Then use Equations 13 and 16 and the property $\left|\sigma_{n}\bigtriangleup\mathcal{P}\right|/n\to 0$ to deduce
$\frac{1}{n}\mathrm{LDS}_{r\sqrt{n}}(\mathcal{P})\geq F_{\mathrm{LSKV}}(r)-5/K$ (20)
@@
Finally, we conclude the proof of Theorem 2.7 with Equations 15 and 20, as they hold for any $K$. ∎
@@
With the same arguments as in the proof of [Sjö23, Lemma 10.1], we could refine our proof of Theorem 2.7 to show that the longest decreasing subsequences of $\mathcal{P}$ are located on the same limit curves as in the uniform case (see [Sjö23, Theorem 10.2 (a)] for a precise statement).
Proof of 2.8.
@@
where $\epsilon_{n}:=\epsilon-\frac{r\sqrt{\check{n}}}{\check{n}}\underset{n\to\infty}{\longrightarrow}\epsilon$ in probability. Then by Theorem 2.7 and dominated convergence theorem:

### v4

@@ 5.2 The uniform case
5.3 Application to our setting
We are now ready to prove Theorem˜2.7. Thanks to ˜3.3, the proof is very close to that of Theorem˜5.8; the main difference is that we need to work individually on a family of rectangular domains which do not intersect the diagonal of $(0,1)^{2}$, and we combine them together with the same techniques as in the proof of Theorem˜5.8.
Proof of Theorem˜2.7.
Fix $r\geq 0$. Let $\mathcal{P}$ be a geometric construction of $\tau_{n}$. As in the proof of Theorem˜5.8, we wish to bound $\left|D(w\sqrt{n})\cap\mathcal{P}\right|$ above and below, uniformly over $w\in\mathcal{U}_{0,r}((0,1)^{2})$. Consider two essentially disjoint intervals $I,J$ of $[0,1]$ and write $\bigtriangleup$ for the symmetric difference of sets. Thanks to ˜3.2 and ˜3.3, one can construct a homogeneous Poisson point process with intensity $n$ on $I\times J$, say $\sigma_{n}$, such that ${\left|\sigma_{n}\bigtriangleup\mathcal{P}\right|}/{n}\to 0$ as $n\to\infty$ almost surely.
@@ 5.3 Application to our setting
Figure 5: The bands defined in the proof of Theorem˜2.7. On each $A_{i}$ and each $B_{i}$, $\mathcal{P}$ is almost distributed like a homogeneous Poisson point process.
@@
for $i\in[K]$, see Figure˜5. These bands each fit in the setting of ˜3.3, and they satisfy $\mathrm{Leb}(C)=\frac{1}{K+1}$ where $C:=(0,1)^{2}\setminus\left(\bigcup_{i}\bar{A}_{i}\cup\bar{B}_{i}\right)$. Moreover, using ˜3.2 and ˜3.3 as above, one can construct a point process $\sigma_{n}$ on $(0,1)^{2}$ such that:
@@
For each $i\in[K]$, we can apply ˜5.2 to $\sigma_{n}$ on $A_{i}$. Hence for each $u\in\mathcal{U}(A_{i})$, there exists $\delta_{u,A_{i}}>0$ such that:
$\sup_{w\in\mathcal{U}(A_{i}),\,\lVert w-u\rVert_{1,A_{i}}<\delta_{u,A_{i}}}\frac{1}{n}\left|D(w\sqrt{n})\cap\sigma_{n}\cap A_{i}\right|<F_{A_{i}}(u)+\epsilon$
@@
$\sup_{w\in\mathcal{U}(C)}\frac{1}{n}\left|D(w\sqrt{n})\cap\sigma_{n}\cap C\right|<\frac{1}{n}\left|\sigma_{n}\cap C\right|=\frac{1}{n}\left|\mathcal{P}\cap C\right|.$
Moreover:
$F_{(0,1)^{2}}(u)=F_{C}(u)+\sum_{i=1}^{K}(F_{A_{i}}(u)+F_{B_{i}}(u))\geq\sum_{i=1}^{K}\left(\big.F_{A_{i}}(u)+F_{B_{i}}(u)\right).$
By summing the previous inequalities we deduce that for each $u\in\mathcal{U}\left((0,1)^{2}\right)$, with $\delta_{u}:=\min_{1\leq i\leq K}\delta_{u,A_{i}}\wedge\delta_{u,B_{i}}$:
$\sup_{w\in\mathcal{U}\left((0,1)^{2}\right),\,\lVert w-u\rVert_{1,(0,1)^{2}}<\delta_{u}}\frac{1}{n}\left|D(w\sqrt{n})\cap\sigma_{n}\right|<F_{(0,1)^{2}}(u)+2K\epsilon+\frac{1}{n}\left|\mathcal{P}\cap C\right|$
@@
$\sup_{w\in\mathcal{U}\left((0,1)^{2}\right),\,\lVert w-u\rVert_{1,(0,1)^{2}}<\delta_{u}}\frac{1}{n}\left|D(w\sqrt{n})\cap\mathcal{P}\right|<F_{(0,1)^{2}}(u)+2K\epsilon+\frac{1}{n}\left|\mathcal{P}\cap C\right|+\frac{1}{n}\left|\sigma_{n}\bigtriangleup\mathcal{P}\right|$
w.h.p. as $n\to\infty$. Recall that $\epsilon=1/K^{2}$ and that $\frac{1}{n}\left|\sigma_{n}\bigtriangleup\mathcal{P}\right|\to 0$ a.s. as $n\to\infty$. Moreover, ˜3.2 implies that $\frac{1}{n}\left|\mathcal{P}\cap C\right|\to\mathrm{Leb}(C)=1/(K+1)$ a.s. as $n\to\infty$. Hence for each $u\in\mathcal{U}\left((0,1)^{2}\right)$:
$\sup_{w\in\mathcal{U}\left((0,1)^{2}\right),\,\lVert w-u\rVert_{1,(0,1)^{2}}<\delta_{u}}\frac{1}{n}\left|D(w\sqrt{n})\cap\mathcal{P}\right|<F_{(0,1)^{2}}(u)+4/K$
w.h.p. as $n\to\infty$. Using the fact that $\mathcal{U}_{0,r}\left((0,1)^{2}\right)$ is compact for the $L^{1}$ metric, as in the proof of Theorem˜5.8, we get:
$\sup_{w\in\mathcal{U}_{0,r}((0,1)^{2})}\frac{1}{n}\left|D(w\sqrt{n})\cap\mathcal{P}\right|<\sup_{u\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)}F_{(0,1)^{2}}(u)+4/K$ (24)
@@
This bound is almost the same as the upper bound of Theorem˜5.8, mutatis mutandis. Fix an arbitrary $u\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)$; we wish to exhibit a well-crafted $w_{n}\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)$ such that $\frac{1}{n}\left|D(w_{n}\sqrt{n})\cap\mathcal{P}\right|\geq F_{(0,1)^{2}}(u)-\epsilon$ holds w.h.p. as $n\to\infty$.
@@
First apply ˜5.5 to each domain $A_{i}\cup B_{i}$ for $i\in[K]$. For each $0<\iota<1/2$, let $S_{\iota}^{i}$ and $\mathfrak{P}_{\iota}^{i}$ be the resulting sets. Then define $S_{\iota}:=\bigcup_{i\in[K]}S_{\iota}^{i}$ and $\mathfrak{P}_{\iota}:=\bigcup_{i\in[K]}\mathfrak{P}_{\iota}^{i}$. Denote by $\widetilde{\mathfrak{P}}_{\iota}$ the set of well-behaved parallelograms in $\mathfrak{P}_{\iota}$. For each $P\in\mathfrak{P}_{\iota}$, let $P^{\prime}$ denote the open parallelogram obtained by shrinking $P$ by a factor $1\!-\!2\iota$ in width and height, as in ˜5.7.
@@
Note that any parallelogram contained in $A_{i}\cup B_{i}$ is either wholly contained in $A_{i}$, or wholly contained in $B_{i}$. Since $\sigma_{n}$ is distributed like a homogeneous Poisson point process with intensity $n$ on each $A_{i}$ and each $B_{i}$, we can use ˜5.3: for each $P\in\widetilde{\mathfrak{P}}_{\iota}$, w.h.p. as $n\to\infty$, there exists $w_{P}^{(n)}\in\mathcal{U}_{u(x_{P},y_{P})-(1-2\iota)\iota c_{P},2(1-2\iota)\iota c_{P}}(P^{\prime})$ such that
@@
$\frac{1}{n\cdot\mathrm{Leb}(P^{\prime})}\left|D\left(w_{P}^{(n)}\sqrt{n}\right)\cap\sigma_{n}\cap P^{\prime}\right|\geq L(u_{x}^{P}u_{y}^{P})-o_{\iota}(1)$ (25)
where $o_{\iota}(1)\to 0$ as $\iota\to 0$, uniformly in $P$ and $n$. This also holds for any $P\in\mathfrak{P}_{\iota}\setminus\widetilde{\mathfrak{P}}_{\iota}$ and any $w_{P}^{(n)}\in\mathcal{U}(P^{\prime})$, for the same reason as in the proof of Theorem˜5.8.
@@
Under the event that $w_{P}^{(n)}$ exists for all $P\in\widetilde{\mathfrak{P}}_{\iota}$, define a function $w_{n}$ on $\bigcup_{P\in\widetilde{\mathfrak{P}}_{\iota}}P^{\prime}\,\cup\,\left((0,1)^{2}\setminus\bigcup_{P\in\widetilde{\mathfrak{P}}_{\iota}}\mathring{P}\right)$ as follows: $w_{n}$ coincides with $w_{P}^{(n)}$ on $P^{\prime}$ for each $P\in\widetilde{\mathfrak{P}}_{\iota}$, whereas is coincides with $u$ on $(0,1)^{2}\setminus\bigcup_{P\in\widetilde{\mathfrak{P}}_{\iota}}\mathring{P}$. As in the proof of Theorem˜5.8, one can check that $w_{n}$ is doubly increasing. Then by ˜5.6, this function can be extended to $w_{n}\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)$. Therefore, using (25), we can write w.h.p. as $n\to\infty$:
@@
$\frac{1}{n}\left|D(w_{n}\sqrt{n})\cap\sigma_{n}\right|\geq\sum_{P\in\mathfrak{P}_{\iota}}\frac{1}{n}\left|D(w_{n}\sqrt{n})\cap\sigma_{n}\cap P^{\prime}\right|\geq\sum_{P\in\mathfrak{P}_{\iota}}\mathrm{Leb}(P^{\prime})\left(L(u_{x}^{P}u_{y}^{P})-o_{\iota}(1)\right)$ (26)
where $o_{\iota}(1)$ is independent of $P$ and $n$. By item (g) of ˜5.5:
$\sum_{P\in\mathfrak{P}_{\iota}}\mathrm{Leb}(P^{\prime})L(u_{x}^{P}u_{y}^{P})=(1\!-\!2\iota)^{2}\sum_{P\in\mathfrak{P}_{\iota}}\mathrm{Leb}(P)L(u_{x}^{P}u_{y}^{P})\geq(1\!-\!2\iota)^{2}F_{\bigcup_{\mathfrak{P}_{\iota}}\!\!P}(u)-o_{\iota}(1).$ (27)
Let us also use [S23, Lemma 4.9], which states that $0\leq L(\partial_{x}u\,\partial_{y}u)\leq 1$. Along with items (a) and (c) of ˜5.5, we get:
$\mathrm{Leb}\left(\big.\bigcup_{\mathfrak{P}_{\iota}}P\right)\geq\mathrm{Leb}\left(\bigcup_{i\in[K]}A_{i}\cup B_{i}\right)-K\epsilon-o_{\iota}(1)=1-2/K-o_{\iota}(1).$
@@
This implies $F_{\bigcup_{\mathfrak{P}_{\iota}}\!\!P}(u)\geq F_{(0,1)^{2}}(u)-2/K-o_{\iota}(1)$. Along with (26) and (27), this yields:
$\frac{1}{n}\left|D(w_{n}\sqrt{n})\cap\sigma_{n}\right|\geq F_{(0,1)^{2}}(u)-2/K-o_{\iota}(1)$
@@
w.h.p. as $n\to\infty$, where $o_{\iota}(1)$ is uniform in $n$. Take $\iota$ such that this $o_{\iota}(1)$ is less than $1/K$. Using the facts that this holds for any $u\in\mathcal{U}_{0,r}\left((0,1)^{2}\right)$ and that $\left|\sigma_{n}\bigtriangleup\mathcal{P}\right|/n\to 0$, we deduce that:
$\sup_{w\in\mathcal{U}_{0,r}((0,1)^{2})}\frac{1}{n}\left|D(w\sqrt{n})\cap\mathcal{P}\right|\geq\sup_{u\in\mathcal{U}_{0,r}((0,1)^{2})}F_{(0,1)^{2}}(u)-5/K$ (28)
@@
Putting Equations˜24 and 28 together, along with (23), we get that:
$\sup_{w\in\mathcal{U}_{0,r}((0,1)^{2})}\frac{1}{n}\left|D(w\sqrt{n})\cap\mathcal{P}\right|\,\longrightarrow\,\sup_{u\in\mathcal{U}_{0,r}((0,1)^{2})}F_{(0,1)^{2}}(u)=F_{\mathrm{LSKV}}(r)$
in probability as $n\to\infty$. Thanks to (13), this concludes the proof. ∎
@@
With the same arguments as in the proof of [S23, Lemma 10.1], we could refine our proof of Theorem˜2.7 to show that the longest decreasing subsequences of $\mathcal{P}$ are located on the same limit curves as in the uniform case (see [S23, Theorem 10.2 (a)] for a precise statement).
Proof of ˜2.8.
@@
where $\epsilon_{n}:=\epsilon-\frac{r\sqrt{\check{n}}}{\check{n}}\underset{n\to\infty}{\longrightarrow}\epsilon$ in probability. Then by Theorem˜2.7 and dominated convergence theorem: